# On five questions of Bourin and Lee: symmetric moduli and an Euler operator identity

### v2

@@ 3.2. A $2\times 2$ counterexample for $p>2$
$Z_{\theta}^{*}Z_{\theta}=\begin{pmatrix}c\\ -s\end{pmatrix}(c,\,-s)=\begin{pmatrix}1&0\\ 0&0\end{pmatrix}=:P.$
@@
Combining (3.1), Claims 3.1 and 3.2, we see that if (1.2) were to hold for $Z_{\theta}$, then necessarily
@@
Assume $p>2$. Consider the $2\times 2$ matrix $Z_{\theta}$ defined in (3.2). If (1.2) held for all $Z\in\mathbb{M}_{2}$, then it would hold for $Z_{\theta}$ and imply the trace inequality (3.1). Claims 3.1 and 3.2 reduce (3.1) to (3.5), which fails for all sufficiently small $\theta$ by Claim 3.3. Hence (1.2) cannot hold for all $Z\in\mathbb{M}_{2}$ when $p>2$.
@@
The construction is intentionally minimal and answers the existence question in Question 1.9. Here $Z$ has rank $2$, and the operator $\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}$ is diagonal and easy to read off explicitly. The strict inequality is obtained by selecting indices $(j,k)=(2,0)$ for which the corresponding singular value of $\operatorname{Re}Z$ vanishes at level $1+j=3$, while $\operatorname{Im}Z$ contributes only its top singular value $2^{-1/2}$, yet the third singular value on the left-hand side equals $2^{-1/p}>2^{-1/2}$ whenever $p>2$.
5. Rectangular Thompson inequality and proof of Theorem 1.12
In this section, we establish a rectangular version of Thompson’s inequality, which will be the main tool in the proof of Theorem 1.12 (ii). The key idea is to reduce the rectangular setting to the square one by means of the polar decomposition: for a rectangular matrix $X\in\mathbb{M}_{m,n}$, one can write $X=W|X|$ with a partial isometry $W$ whose initial projection coincides with the support projection of $|X|$. After recording this elementary fact (Lemma 5.1), we apply the classical (square) Thompson inequality to the compressed matrices $W^{*}A$ and $W^{*}B$ and then compare $|W^{*}A|,|W^{*}B|$ with $|A|,|B|$, respectively, to obtain the desired unitary-orbit domination. Finally, we use the rectangular Thompson inequality to settle Question 1.11, thereby proving Theorem 1.12.
Lemma 5.1.
Let $X\in\mathbb{M}_{m,n}$. Then there exists a partial isometry $W\in\mathbb{M}_{m,n}$ such that
$X=W|X|,\qquad|X|=(X^{*}X)^{1/2},$
and
$W^{*}W=\mathrm{supp}(|X|),\qquad W^{*}W\leq I_{n},WW^{*}\leq I_{m},\qquad(W^{*}W)\,|X|=|X|,$
where $\mathrm{supp}(\cdot)$ means the support projection.
Proof.
Take a singular value decomposition $X=U\Sigma V^{*}$, where $U\in\mathbb{M}_{m}$ and $V\in\mathbb{M}_{n}$ are unitary and
$\Sigma=\begin{pmatrix}D&0\\ 0&0\end{pmatrix}\in\mathbb{M}_{m,n},\qquad D=\operatorname{diag}(s_{1},\dots,s_{r}),\ \ s_{1}\geq\cdots\geq s_{r}>0,\ \ r=\operatorname{rank}(X).$
Then
$X^{*}X=V(\Sigma^{*}\Sigma)V^{*},\qquad\Sigma^{*}\Sigma=\begin{pmatrix}D^{2}&0\\ 0&0\end{pmatrix}$,
hence
$|X|=(X^{*}X)^{1/2}=V\begin{pmatrix}D&0\\ 0&0\end{pmatrix}V^{*}.$
$W|X|=U\begin{pmatrix}I_{r}&0\\ 0&0\end{pmatrix}V^{*}\cdot V\begin{pmatrix}D&0\\ 0&0\end{pmatrix}V^{*}=U\begin{pmatrix}D&0\\ 0&0\end{pmatrix}V^{*}=U\Sigma V^{*}=X.$
Moreover,
$(W^{*}W)|X|=P|X|=\Bigl(V\begin{pmatrix}I_{r}&0\\ 0&0\end{pmatrix}V^{*}\Bigr)\Bigl(V\begin{pmatrix}D&0\\ 0&0\end{pmatrix}V^{*}\Bigr)=V\begin{pmatrix}D&0\\ 0&0\end{pmatrix}V^{*}=|X|.$
That is, $W^{*}W$ is exactly the support projection of $|X|$. ∎
Lemma 5.1 allows us to choose the partial isometry in the polar decomposition so that it acts as the identity on the range of $|X|$. This normalization is exactly what we need to turn $|A+B|$ into the absolute value of a square matrix, to which the classical Thompson inequality applies.
Lemma 5.2 (Rectangular Thompson inequality).
Let $A,B\in\mathbb{M}_{m,n}$. Then there exist unitary matrices $U,V\in\mathbb{M}_{n}$ such that
$|A+B|\ \leq\ U|A|U^{*}+V|B|V^{*}.$ (5.1)
Proof.
$W^{*}(A+B)=W^{*}\left(W|A+B|\right)=|A+B|.$
In particular,
$|A+B|=|W^{*}(A+B)|=|W^{*}A+W^{*}B|.$ (5.2)
Now $W^{*}A$ and $W^{*}B$ are $n\times n$ (square) matrices. Applying Thompson’s inequality (1.1) (the square case) to $W^{*}A$ and $W^{*}B$, there exist unitaries $U,V\in\mathbb{M}_{n}$ such that
$|W^{*}A+W^{*}B|\leq U\,|W^{*}A|\,U^{*}+V\,|W^{*}B|\,V^{*}.$ (5.3)
Since $WW^{*}\leq I_{m}$, we have
$|W^{*}A|^{2}=(W^{*}A)^{*}(W^{*}A)=A^{*}WW^{*}A\leq A^{*}A=|A|^{2}.$
$U|W^{*}A|U^{*}\leq U|A|U^{*},\qquad V|W^{*}B|V^{*}\leq V|B|V^{*}.$
Substituting into (5.3) and combining with (5.2) yields (5.1). ∎
With Lemma 5.2 in hand, we now prove Theorem 1.12. Part (i) is verified by a concrete counterexample for the symmetric modulus, while part (ii) follows from an application of Lemma 5.2 to a suitable linear embedding that converts the quadratic symmetrization into a rectangular modulus.
Proof of Theorem 1.12.
(i) Take
$X=\begin{pmatrix}-1&-1\\ 0&-1\end{pmatrix},\qquad Y=\begin{pmatrix}0&-1\\ 0&0\end{pmatrix}\in M_{2}(\mathbb{R}).$
A direct computation gives
$|X|_{\mathrm{sym}}=\frac{1}{2\sqrt{5}}\begin{pmatrix}5&2\\ 2&5\end{pmatrix},\qquad|Y|_{\mathrm{sym}}=\frac{1}{2}I_{2},\qquad|X+Y|_{\mathrm{sym}}=\frac{1}{\sqrt{2}}\begin{pmatrix}2&1\\ 1&2\end{pmatrix}$.
Hence
$\bigl\||X+Y|_{\mathrm{sym}}\bigr\|_{\infty}=\frac{3}{\sqrt{2}},\qquad\bigl\||X|_{\mathrm{sym}}\bigr\|_{\infty}=\frac{7}{2\sqrt{5}},\qquad\bigl\||Y|_{\mathrm{sym}}\bigr\|_{\infty}=\frac{1}{2},$
so
$\bigl\||X+Y|_{\mathrm{sym}}\bigr\|_{\infty}>\bigl\||X|_{\mathrm{sym}}\bigr\|_{\infty}+\bigl\||Y|_{\mathrm{sym}}\bigr\|_{\infty}\qquad\left(\text{i.e. }\ \frac{3}{\sqrt{2}}>\frac{7}{2\sqrt{5}}+\frac{1}{2}\right).$
If the inequality in Question 1.11 held for some unitaries $U,V$, then taking operator norms would yield
$\bigl\||X+Y|_{\mathrm{sym}}\bigr\|_{\infty}\leq\bigl\|U|X|_{\mathrm{sym}}U^{*}+V|Y|_{\mathrm{sym}}V^{*}\bigr\|_{\infty}\leq\bigl\||X|_{\mathrm{sym}}\bigr\|_{\infty}+\bigl\||Y|_{\mathrm{sym}}\bigr\|_{\infty},$
a contradiction. This proves (i).
(ii) Define the linear map $\Gamma:\mathbb{M}_{n}\to\mathbb{M}_{2n,n}$ by
$\Gamma(Z):=\begin{pmatrix}Z\\ Z^{*}\end{pmatrix}$.
Then
$\Gamma(Z)^{*}\Gamma(Z)=Z^{*}Z+ZZ^{*},\qquad\text{hence}\qquad|\Gamma(Z)|=\bigl(Z^{*}Z+ZZ^{*}\bigr)^{1/2}=\sqrt{2}\,|Z|_{\mathrm{qsym}}.$
Moreover, $\Gamma(X+Y)=\Gamma(X)+\Gamma(Y)$. Applying rectangular Thompson inequality (Lemma 5.2) yields unitaries $U,V\in\mathbb{M}_{n}$ such that
$|\Gamma(X+Y)|\leq U|\Gamma(X)|U^{*}+V|\Gamma(Y)|V^{*}.$
Dividing by $\sqrt{2}$ and using $|\Gamma(Z)|=\sqrt{2}\,|Z|_{\mathrm{qsym}}$ gives
$|X+Y|_{\mathrm{qsym}}\leq U|X|_{\mathrm{qsym}}U^{*}+V|Y|_{\mathrm{qsym}}V^{*},$
as desired. ∎

### v3

@@ 9.2. A $2\times 2$ counterexample for $p>2$
$Z_{\theta}^{*}Z_{\theta}=\begin{pmatrix}c&-s\\ 0&0\end{pmatrix}\begin{pmatrix}c&0\\ -s&0\end{pmatrix}=\begin{pmatrix}1&0\\ 0&0\end{pmatrix}=:P.$
@@
Combining (9.1), Claims 9.1 and 9.2, we see that if (1.10) were to hold for $Z_{\theta}$, then necessarily
@@
Assume $p>2$. Consider the $2\times 2$ matrix $Z_{\theta}$ defined in (9.2). If (1.10) held for all $Z\in\mathbb{M}_{2}$, then it would hold for $Z_{\theta}$ and imply the trace inequality (9.1). Claims 9.1 and 9.2 reduce (9.1) to (9.5), which fails for all sufficiently small $\theta$ by Claim 9.3. Hence (1.10) cannot hold for all $Z\in\mathbb{M}_{2}$ when $p>2$.
@@
The construction is intentionally minimal and answers the existence question in Question 1.33. Here $Z$ has rank $2$, and the operator $\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}$ is diagonal and easy to read off explicitly. The strict inequality is obtained by selecting indices $(j,k)=(2,0)$ for which the corresponding singular value of $\operatorname{Re}Z$ vanishes at level $1+j=3$, while $\operatorname{Im}Z$ contributes only its top singular value $2^{-1/2}$, yet the third singular value on the left-hand side equals $2^{-1/p}>2^{-1/2}$ whenever $p>2$.